OpenStudy (barreraa):

Please Help !!!!!!!!!!!!!!!!!!!!!!!Find the solution set for the open sentence with the given replacement set. w2 + 16 = 8w, {0, 2, 4, 8} (Points : 1) 0 2 4 8

OpenStudy (freckles):

It sounds like you are most to use the given set to find the solution.

OpenStudy (freckles):

try putting in the numbers one at a time and see if they work

OpenStudy (barreraa):

ok

OpenStudy (freckles):

for example if we replace w with 0 we have 0^2+16=8(0) <--is this true?

OpenStudy (freckles):

by "is this true?" I'm asking are both sides the same

OpenStudy (barreraa):

yes

OpenStudy (freckles):

0^2=0*0=0 and 0(8)=0 so you are saying 16=0?

OpenStudy (barreraa):

oh no

OpenStudy (freckles):

oh no is right :)

OpenStudy (freckles):

now try pluggin in the second value in your set

OpenStudy (freckles):

replace the w's with 2's now

OpenStudy (freckles):

2^2+16=8(2)

OpenStudy (freckles):

are both sides the same?

OpenStudy (barreraa):

no

OpenStudy (freckles):

right because 2^2+16=4+16=20 and 20 is definitely not 8(2)=16

OpenStudy (freckles):

so we ruled out so far that w=0 or w=2 are solutions

OpenStudy (freckles):

let's try w=4

OpenStudy (freckles):

replace the w's with 4's

OpenStudy (freckles):

4^2+16=8(4) <---is this true?

OpenStudy (barreraa):

no

OpenStudy (freckles):

remember 4^2 means 4*4

OpenStudy (freckles):

what is 4*4+16=?

OpenStudy (barreraa):

32

OpenStudy (freckles):

and what is 8(4)

OpenStudy (barreraa):

32

OpenStudy (freckles):

32 is 32

OpenStudy (freckles):

therefore 4^2+16=8(4) is true

OpenStudy (barreraa):

ok thanks so much

OpenStudy (freckles):

np